Find the critical value of x2 based on the given information H1:0< 1000
629
n = 20
a = 0.025

H1:0 629 1000 n = 20 a = 0.025 8.231 32.852 8.907 9.591
Answer #1

Here it is left tailed test.(less than type)

So we have a chi-square table for the right-tailed test.

alpha = 0.025

n= 20

df = 20-1 = 19

X2df,1-alpha = X219,1-0.025 =   X219,0.975 = 8.907

The critical value = 8.907

Answer :- 8.907